# From Propeller Damage Estimation and Adaptation to Fault Tolerant Control: Enhancing Quadrotor Resilience

### v1

@@ III-C1 State Predictor
$\mathbf{\hat{z}}[k]=\mathbf{z}[k-1]+dt\cdot\begin{bmatrix}\dot{\mathbf{v}}[k-1]\\ \dot{\mathbf{\Omega}}[k-1]\end{bmatrix},$ (11)
where $\mathbf{\hat{z}}[k]$ is the predicted state for time $k$. Let the true observed state $\mathbf{z}$ be the values observed from our state estimation. We set a diagonal gain defined by the user $\mathbf{K}\in\mathbb{R}^{6\times 6}$. The state transition is
$\mathbf{\dot{v}}=g\mathbf{e}_{3}+\frac{(f+f_{L1})}{m}\mathbf{R}\mathbf{e}_{3}+\frac{1}{m}\mathbf{R}\bm{\sigma}_{1:3}+\mathbf{K_{1:3}}(\mathbf{\hat{z}}_{1:3}-\mathbf{z}_{1:3}),$ (12)
@@
$\mathbf{\dot{\Omega}}=\mathbf{J}^{-1}(\mathbf{M}+\mathbf{M}_{L1}+\sigma_{4:6}-\mathbf{\Omega}\times\mathbf{J}\mathbf{\Omega})+\mathbf{K_{4:6}}(\mathbf{\hat{z}}_{4:6}-\mathbf{z}_{4:6}).$ (13)
The terms $f_{L1}$ and $\mathbf{M}_{L1}$ are the additional force and moment commanded the L1 adaptive controller.
@@
  Max Speed [$\mathrm{m}\mathrm{/}\mathrm{s}$]$ L1 OFF L1 ON L1 OFF L1 ON L1 OFF L1 ON L1 OFF L1 ON L1 off
12 $[0.5]$ $x$ 0.042 0.031 0.072 0.017 0.127 0.020 - 0.207 0.041
$y$ 0.092 0.070 0.233 0.062 0.453 0.030 - 0.358 0.044
$z$ 0.017 0.002 0.050 0.002 0.082 0.003 - 0.005 0.011
8 $[0.8]$ $x$ 0.040 0.037 0.067 0.031 0.16 0.032 - 0.208 0.075
$y$ 0.10 0.064 0.222 0.064 0.42 0.048 - 0.384 0.059
$z$ 0.022 0.002 0.051 0.004 0.09 0.003 - 0.007 0.020
5 $[1.1]$ $x$ 0.051 0.045 0.079 0.060 0.181 0.046 - 0.204 0.093
$y$ 0.112 0.079 0.249 0.072 0.406 0.047 - 0.406 0.109
$z$ 0.025 0.003 0.050 0.005 0.090 0.003 - 0.009 0.019
III-C2 L1 Adaptation Law
We can characterize the external disturbances on our force and moments. We use the previously selected $\mathbf{K}$ diagonal matrix to generates a specific gain, $\bm{\phi}$, to drive the disturbance exponentially to zero
$\begin{split}\bm{\phi}&=(\exp(\mathbf{K}\cdot dt-\mathbf{I})^{-1}\mathbf{K})\exp(\mathbf{K}\cdot dt),\\ \bm{\sigma}&=\begin{bmatrix}m\mathbf{R}^{\top}&0\\ 0&\mathbf{J}\\ \end{bmatrix}\bm{\phi}(\mathbf{z}[k]-\mathbf{\hat{z}}[k]).\end{split}$ (14)
We then estimate our acceleration and moment disturbances using eq. (14) through mapping the velocity and angular velocity errors between our state predictor and estimation to calculate disturbances $\bm{\sigma}$. We solve the additional action $\bm{\mu}_{L1}=\begin{bmatrix}f_{L1}&\mathbf{M}_{L1}^{\top}\end{bmatrix}^{\top}$ from our disturbance $\bm{\sigma}$. The moment disturbance estimates $\bm{\sigma}_{4:6}$ and forces acting along the thrust axis $\bm{\sigma}_{3}$ can be negated by providing an equal but negative compensation to cancel the disturbance $\bm{\mu}_{L1}=-\bm{\sigma}_{3:6}$. We also implement an exponential Low Pass Filter (LPF) to effectively mitigate the noise present in the state estimation, thereby achieving smoother results: $\bm{\mu}_{L1}=\leavevmode\nobreak\ -\text{LPF}(\bm{\sigma}_{3:6})$.
@@ III-D Propeller Damage Estimation
We estimate the propellers’ damage using our augmented control. We assume that in nominal flight conditions a quadratic relationship between the $i^{\text{th}}$ propeller thrust and corresponding motor speed $f_{i}=k_{f_{\text{real}}}\omega_{i}^{2}$ where $k_{f_{\text{real}}}$ represents the true thrust coefficient of the propeller and $\omega_{i}$ is the motor speed. A damaged propeller spinning at the original RPM provides lower thrust due to the change in $k_{f_{\text{real}}}$. We quantify our propeller damages on the thrust coefficient by
@@
This threshold is set as undamaged propellers are likely to have negative damage due to torque mismatch. However a negative damage is physically impossible to exist. This means all undamaged propellers will have negative or close to zero damage. A detailed explanation of this concept is discussed in Section V.
@@
Figure 4: 3D circular trajectory tracking trace for L1 Off (left) and L1 On (right). Each color of trajectory represents a different percentage of propeller damage as stated in the legend. The plot slows that the position drift grows as the kf mismatch gets larger. The plot on the right shows that the L1 adaptive algorithm is able to compensate for such propeller damages.
Figure 5: RMSE averaged over 3 axes vs Propeller Damage at hover with various levels of propeller damage.
@@
where $\mathbf{b}$ represents the desired force and moments of the quadrotor commanded by the nominal control without L1 and $\bm{\omega}_{\text{L1}}=\begin{bmatrix}\omega_{{L1}_{1}}&\omega_{{L1}_{2}}&\omega_{{L1}_{3}}&\omega_{{L1}_{4}}\end{bmatrix}^{\top}$ are the RPMs from the combination of L1 and nominal geometric system, $d_{x}$ and $d_{y}$ represent the distance from the propeller to the center of mass in the body frame projected on axes $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ respectively. To prevent the torque effect from L1 control we removed the fourth row of our constraints along with $M_{3}$ (the torque components) from our optimization. In addition, torque coefficients are difficult to observe and calculate without inducing yaw movement in the quadrotor. The L1 controller drives the error to zero between desired and executed action through the use of supplemental control action[21]. We construct our equality constraint in eq. (20) to represent this behaviour where the real thrust coefficients must be consistent with the supplemental action that generates the desired action.
The optimum based on null-space projection is
@@ III-E Fault-Tolerant Transition
We leverage our propeller damage estimate, eq. (21) to determine when to switch to fault-tolerant control. Our fault-tolerant control is implemented based on our previous work [13] where the approach is only tested in simulation. From the qualitative results with damaged propellers, we set a threshold at $50\%$ damage. Based on our vehicle’s thrust-to-weight ratio of $2.5:1$, we experimentally verify that $50\%$ damage provides a safety margin within the limit. A rapid response is necessary because a quadrotor could fall several meters if the estimation has a slow rise time. This is achieved by careful tuning of our LPF to balance denoising and response time. Once the threshold is met, the fault-tolerant control is transitioned from the base per our control design Fig. 3.
Figure 7: Simultaneous propeller damage (propellers 2 and 3) for MAV moving $0.5$\mathrm{m}\mathrm{/}\mathrm{s}$$ in a circle
Figure 8: Damage injected to propeller $1$. Fault-tolerant transition threshold is set to $50\%$. Real-world experiment starts with $40\%$ damage and transitions when $50\%$ damage is injected. These experiments correspond to each row in Fig. (9).
@@ IV Results
We conduct our experiments in a flying space of $10\times 6\times 4\leavevmode\nobreak\ $\mathrm{m}^{3}$$ at the Agile Robotics and Perception Lab (ARPL) at New York University to validate the proposed approaches. The environment is equipped with a Vicon motion capture system which provides accurate pose estimates at $100\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. This is fused with IMU measurements through an Unscented Kalman Filter to provide state estimates at $500\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. The robot, based on our previous work [23], is equipped with a VOXL®2 ModalAITM and four brushless motors and modified to obtain a $2.5$ to $1$ thrust to weight ratio with total a weight of $700\leavevmode\nobreak\ $\mathrm{g}$$. We classify the $4$ levels of propeller damage based on $k_{f_{\text{mis}}}$: No damage refers to $k_{f_{\text{mis}}}=0\%$. Low damage refers to $k_{f_{\text{mis}}}=20\%$. Medium Damage refers to $k_{f_{\text{mis}}}=40\%$. High damage refers to $k_{f_{\text{mis}}}=60\%$.
Figure 9: Position tracking performance after propeller damage is injected. Each row corresponds to the propeller damage estimate results in Fig. 8. Top row: $30\%$ prop damage is injected so no fault-tolerant control is activated. Middle row: $80\%$ is injected triggering fault-tolerant control. Bottom row: $50\%$ is injected in real world flight triggering fault control.
@@ IV-A L1 Performance Study
In order to evaluate the overall performance, the quadrotor is commanded to follow an ellipse of radius $1$ m in $x$, $0.6$ m in $y$, and $0.1$ m in $z$ for each case of propeller damage as well as a fully failed single propeller case using fault-tolerant control. The trajectory is completed with and without L1 adaptive control. This ellipse is flown at three different speeds: $12$, $8$, and $5$ s periods. As shown in Table I, we observe that the RMSE reduces when the L1 adaptation is switched on in all cases and has lower tracking error than fault-tolerant for propeller damages below $40\%$. Furthermore, in Fig. 4, we depict the $8$ s circular trajectory in $3$D space showing that L1 adaptation compensates better to various propeller damages than with no L1 adaptation. We also see the position error grows in the L1 Off case as the severity of propeller damage increases while L1 maintains accurate tracking performance despite the degree of propeller damage. We also observe based on Fig. 5 that when the damage reaches greater than $40\%$, the error starts becoming too prominent for the adaptive control to negate. Therefore, it would be preferable to switch to a fault-tolerant method.
@@ IV-B Propeller Damage Estimation
The L1 adaptive controller has demonstrated suitability and robustness to propeller damage compensation. In this section, we show that eq. (15) is an accurate estimate of the propeller damage ($k_{f_{\text{mis}}}$) even when the aerial robot is in motion. In Fig. 6, we report the estimates of a single damaged propeller while the quadrotor is executing an ellipse trajectory with a max speed of $1.1\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$. This method is able to estimate the error with an error range of $4\%$. In terms of propeller coefficient $4\%$ is a small absolute difference of $3.8\cdot 10^{-10}$ or $8.3\leavevmode\nobreak\ $\mathrm{g}$$ of thrust difference on a $700\leavevmode\nobreak\ $\mathrm{g}$$ drone. Finally, we estimate dual propeller damages simultaneously in Fig. 7. When conducting simultaneous damage estimates, the damaged propellers, specifically $2$ and $3$, align closely with the true propeller damage.

### v2

@@ III-C1 L1 Adaptation Law
$\begin{split}\mathbf{A}&=(\exp(\mathbf{K}\cdot dt-\mathbf{I})^{-1}\mathbf{K})\exp(\mathbf{K}\cdot dt),\\ \bm{\sigma}[k]&=\begin{bmatrix}m\mathbf{R}^{\top}&0\\ 0&\mathbf{J}\\ \end{bmatrix}\mathbf{A}\left(\begin{bmatrix}\mathbf{v}_{m}[k]\\ \mathbf{\Omega}_{m}[k]\end{bmatrix}-\begin{bmatrix}\mathbf{v}_{p}[k]\\ \mathbf{\Omega}_{p}[k]\end{bmatrix}\right).\end{split}$ (11)
We solve the additional action $\bm{\mu}_{L1}=\begin{bmatrix}f_{L1}&\mathbf{M}_{L1}^{\top}\end{bmatrix}^{\top}$ from our disturbance $\bm{\sigma}$. The moment disturbance estimates $\bm{\sigma}_{4:6}$ and disturbance along the thrust axis $\bm{\sigma}_{3}$ can be negated by providing an equal but negative compensation to cancel the disturbance $\bm{\mu}_{L1}=-\bm{\sigma}_{3:6}$. We implement an exponential Low Pass Filter (LPF) to mitigate the noise present in the state estimation, achieving smoother results: $\bm{\mu}_{L1}=\leavevmode\nobreak\ -\text{LPF}(\bm{\sigma}_{3:6})$. $f_{L1}$ and $\mathbf{M}_{L1}$ are added to eq. (5) and eq. (10) respectively and reflected in Fig. (3).
III-C2 State Predictor
Separately, once we solve for the action $\bm{\mu}_{L1}$ and external disturbances $\bm{\sigma}$, we solve for the next predicted velocities of our system, $(\mathbf{v}_{p}[k],\mathbf{\Omega}_{p}[k])\rightarrow(\mathbf{v}_{p}[k+1],\mathbf{\Omega}_{p}[k+1])$. First, we solve for the linear and angular acceleration $\dot{\mathbf{v}}_{p}[k]$ and $\dot{\mathbf{\Omega}}_{p}[k]$ as
$\mathbf{\dot{v}}_{p}[k]=\underbrace{\frac{(f[k]+f_{L1}[k])}{m}\mathbf{R}\mathbf{e}_{3}}_{\text{actuation}}+\underbrace{\frac{1}{m}\mathbf{R}\bm{\sigma}_{1:3}[k]}_{\text{disturbance}}+\underbrace{g\mathbf{e}_{3}}_{\text{gravity}},$ (12)
@@ III-C2 State Predictor
$\mathbf{\dot{\Omega}}_{p}[k]=\mathbf{J}^{-1}(\underbrace{\mathbf{M}[k]+\mathbf{M}_{L1}[k]}_{\text{actuation}}+\underbrace{\bm{\sigma}_{4:6}[k]}_{\text{disturbance}}-\underbrace{\mathbf{\Omega}_{m}[k]\times\mathbf{J}\mathbf{\Omega}_{m}[k]}_{\text{cross}}).$ (13)
Finally, we integrate over a time step $dt$ with a weighted average of the measurement and the predicted velocities. A weighted average is used to filter the noise inherent in the measured velocity
$\begin{bmatrix}\mathbf{v}_{p}[k+1]\\ \mathbf{\Omega}_{p}[k+1]\end{bmatrix}=(1-\lambda)\begin{bmatrix}\mathbf{v}_{p}[k]\\ \mathbf{\Omega}_{p}[k]\end{bmatrix}+\lambda\begin{bmatrix}\mathbf{v}_{m}[k]\\ \mathbf{\Omega}_{m}[k]\end{bmatrix}+dt\begin{bmatrix}\dot{\mathbf{v}}_{p}[k]\\ \dot{\mathbf{\Omega}}_{p}[k]\end{bmatrix}.$ (14)
The weighted average term is $\lambda=\mathbf{K}\cdot dt$. This is predicted value is looped backed into our L1 control law eq. (11) in the next time step $k+1$
@@
  Max Speed [$\mathrm{m}\mathrm{/}\mathrm{s}$]$ L1 ON L1 OFF L1 ON L1 OFF L1 ON L1 OFF L1 ON L1 OFF L1 ON
12 $[0.5]$ $x$ $\mathbf{0.031}$ 0.042 $\mathbf{0.017}$ 0.072 $\mathbf{0.020}$ 0.127 $\mathbf{0.207}$ 0.041
$y$ $\mathbf{0.070}$ 0.092 $\mathbf{0.062}$ 0.233 $\mathbf{0.030}$ 0.453 $\mathbf{0.358}$ 0.044
$z$ $\mathbf{0.002}$ 0.017 $\mathbf{0.002}$ 0.050 $\mathbf{0.003}$ 0.082 $\mathbf{0.005}$ 0.011
8 $[0.8]$ $x$ $\mathbf{0.037}$ 0.040 $\mathbf{0.031}$ 0.067 $\mathbf{0.032}$ 0.16 $\mathbf{0.208}$ 0.075
$y$ $\mathbf{0.064}$ 0.10 $\mathbf{0.064}$ 0.222 $\mathbf{0.048}$ 0.42 $\mathbf{0.384}$ 0.059
$z$ $\mathbf{0.002}$ 0.022 $\mathbf{0.004}$ 0.051 $\mathbf{0.003}$ 0.09 $\mathbf{0.007}$ 0.020
5 $[1.1]$ $x$ $\mathbf{0.045}$ 0.052 $\mathbf{0.060}$ 0.079 $\mathbf{0.046}$ 0.181 $\mathbf{0.204}$ 0.093
$y$ $\mathbf{0.079}$ 0.112 $\mathbf{0.072}$ 0.249 $\mathbf{0.047}$ 0.406 $\mathbf{0.406}$ 0.109
$z$ $\mathbf{0.003}$ 0.025 $\mathbf{0.005}$ 0.050 $\mathbf{0.003}$ 0.090 $\mathbf{0.009}$ 0.019
@@ III-D Propeller Damage Estimation
We estimate the propellers’ damage using our augmented control. We assume that in nominal flight conditions a quadratic relationship between the $i^{\text{th}}$ propeller thrust and corresponding motor speed $f_{i}=k_{f_{\text{real}}}\omega_{i}^{2}$ where $k_{f_{\text{real}}}$ represents the true thrust coefficient of the propeller and $\omega_{i}$ is the motor speed of the $i^{\text{th}}$ propeller. A damaged propeller spinning at the original RPM provides lower thrust due to the change in $k_{f_{\text{real}}}$. We quantify our propeller damages on the thrust coefficient by the following mismatch index
@@
This threshold is set as undamaged propellers are likely to have negative damage due to torque mismatch. However a negative damage is physically impossible. A detailed explanation of this concept is discussed in Section V.
@@
Figure 5: Tracking RMSE averaged over 3 axes vs Propeller Damage at hover with various levels of propeller damage.
@@
where $\mathbf{b}$ represents the desired force and moments of the quadrotor commanded by the nominal control without L1 and $\bm{\omega}_{\text{L1}}=\begin{bmatrix}\omega_{{L1}_{1}}&\omega_{{L1}_{2}}&\omega_{{L1}_{3}}&\omega_{{L1}_{4}}\end{bmatrix}^{\top}$ are the RPMs from the combination of L1 and nominal geometric system, $d_{x}$ and $d_{y}$ represent the distance from the propeller to the center of mass in the body frame projected on axes $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ respectively. To prevent the torque effect from L1 control we remove the fourth row of our constraints along with $M_{3}$ (the torque components) from our optimization. The L1 controller drives the error to zero between desired and executed action through the use of supplemental control action[24]. We construct our equality constraint in eq. (20) to represent this behaviour where the real thrust coefficients must be consistent with the supplemental action that generates the desired action. The optimum based on null-space projection is
@@ III-E Fault-Tolerant Transition
We leverage our propeller damage estimate, eq. (21) to determine when to switch to fault-tolerant control. Our fault-tolerant control is implemented based on our previous work [21] which was only tested in simulation. From the qualitative results with damaged propellers, we set a threshold at $50\%$ damage. Based on our vehicle’s thrust-to-weight ratio of $2.5:1$, we experimentally verify that $50\%$ damage provides a safety margin within the limit. Once the threshold is met, the system automatically transitions to the fault-tolerant control per our control design as shown in Fig. 3.
Figure 6: Simultaneous propeller damage (propellers 2 and 3) for MAV moving $0.5$\mathrm{m}\mathrm{/}\mathrm{s}$$ in a circle. Both PID (green) and L1 (blue) approaches for propeller damage estimation are shown.
Figure 7: Propeller damage estimate of an undamaged vehicle under 3 different experimental conditions. Hover Baseline refers to a hover with no external forces. Figure Eight refers to flying a trajectory in the shape of an eight with a max velocity of $3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$. Hover fan on refers to stabilizing the vehicle under a fan blowing wind at $3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$
@@ IV Results
Our experiments are conducted in a flying space of $10\times 6\times 4\leavevmode\nobreak\ $\mathrm{m}^{3}$$ at the Agile Robotics and Perception Lab (ARPL) at New York University. The environment is equipped with a Vicon motion capture system which provides accurate pose estimates at $100\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. This is fused with IMU measurements through an Unscented Kalman Filter to provide state estimates at $500\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. The robot, based on our previous work [26], is equipped with a VOXL®2 ModalAITM and four brushless motors and modified to obtain a $2.5$ to $1$ thrust to weight ratio with total a weight of $700\leavevmode\nobreak\ $\mathrm{g}$$. We set the $\mathbf{K}$ such that $\lambda=\text{Diag}\begin{bmatrix}0.4&0.4&0.4&0.1&0.1&0.1\end{bmatrix}$.
@@ IV-A L1 Performance Study
To evaluate the tracking performance, the quadrotor is commanded to follow an ellipse of radius $1$ m in $x$, $0.6$ m in $y$, and $0.1$ m in $z$ for each case of propeller damage as well as under fault-tolerant control with one disabled propeller. The trajectory is completed with and without L1 adaptive control. This ellipse is flown at three different speeds: $12$, $8$, and $5\leavevmode\nobreak\ $\mathrm{s}$$ periods. We observe in Table I that the RMSE reduces when the L1 adaptation is switched on in all cases and has lower tracking error than fault-tolerant for propeller damages below $40\%$. We also see in Fig. 5 the position error grows in the L1 Off case as the severity of propeller damage increases while L1 maintains accurate tracking performance despite the degree of propeller damage. However when the damage reaches greater than $40\%$, the error starts becoming too prominent for the adaptive control to negate. In these cases, transitioning to a fault-tolerant controller is preferable.
Figure 8: Real-world experiments with damage injected to propeller $1$. Fault-tolerant transition threshold is set to $50\%$.
Figure 9: Position tracking performance after propeller damage is injected in real-world flight. The $80\%$ damage case shows a high yaw rate because it triggers fault-tolerant control, causing a rapid spin unlike the $30\%$ damage case.
@@ IV-B Propeller Damage Estimation
The L1 adaptive controller has demonstrated suitability and robustness to propeller damage compensation. In this section, we show that Sec. III-D represents an accurate estimate of the propeller damage, $k_{f_{\text{mis}}}$, even when the aerial robot is in motion. To calculate the various levels of propeller damage, we use a thrust bench to calculate the true thrust coefficient of the damaged propeller, $k_{f_{\text{real}}}$ and undamaged propeller, $k_{f_{\text{model}}}$, coefficient. We then use eq. (15) to calculate the damage, $k_{f_{\text{mis}}}$. We test $4$ levels of propeller damage based on $k_{f_{\text{mis}}}$: No damage refers to $k_{f_{\text{mis}}}=0\%$. Low damage refers to $k_{f_{\text{mis}}}=20\%$. Medium Damage refers to $k_{f_{\text{mis}}}=40\%$. High damage refers to $k_{f_{\text{mis}}}=60\%$. In Fig. 4, we report the estimates of a single damaged propeller at the aforementioned levels while the quadrotor is executing an ellipse trajectory with a max speed of $1.1\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$. This method is able to estimate the error with an error range of $4\%$. In terms of individual actuator forces, this represents $8.3\leavevmode\nobreak\ $\mathrm{g}$$ of thrust on a $700\leavevmode\nobreak\ $\mathrm{g}$$ drone. Next, we estimate dual propeller damage simultaneously in Fig. 6. We compare two methods for damage estimation. First, we use L1 complimentary actions shown in blue. Next, we replace the L1 adaptive control with an integral term in a PID controller for complimentary action in green. We substitute $\omega_{L1i}$ with the actions generated by an additional integral controller to estimate the propeller damage following the same inference and optimization procedure as Section III-D. The integral term is placed over the velocity expressed in the body frame. While the integral term can compensate some damage, it is unsuitable for damage estimation. The L1 complimentary actions on the other hand can be used to accurately estimate the propeller’s damage.
Furthermore, we test the effects of medium speed flight ($3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$) and external disturbances on our propeller damage estimation. In Fig. 7, we show the results of 3 experiments. First, we show a control case where the quadrotor is in hover without external disturbances and no damages. Second, we show the same vehicle in hover with $3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$ wind disturbance. Third, we show a flight without external disturbances flying at $3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$. Ideally, the desired estimated damage should be $0\%$. Overall, we notice that the fan has a significant effect on damage estimation around $20\%$, but it is not enough to trigger a transition. Medium speed motions at $3\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$ instead cause minimal but noticeable deviations from around $0\%$ to ±$\pm 10\%$.